# From See to Shield: ML-Assisted Fine-Grained Access Control for Visual Data

### v1

@@ 1. Introduction
The rapid growth of data-driven digital systems and applications has led to the collection and storage of large volumes of sensitive data, ranging from personally identifiable information (PII) to financial transactions and healthcare records. Managing and regulating access to this information has become a critical challenge for individuals and organizations. Beyond protection against external threats, effective access control is a cornerstone of data security, ensuring that only authorized users, applications, and processes can interact with sensitive information. Access control mechanisms regulate and restrict access to resources, systems, or physical areas, typically granting permissions to users based on predefined roles and policies. For example, access to medical records is limited to healthcare professionals, surveillance footage is restricted to security personnel, and digital copies of sealed judicial records are provided only to lawyers and judges. While dealing with such semantically nuanced restrictions, access control mechanisms must operate both reliably and efficiently. The consequences of inadequate access control mechanisms can result in serious financial, reputational, and operational damage as seen in major data breach incidents [^1] https://www.ibm.com/reports/data-breach. Complementing this reality, regulatory frameworks (e.g., the GDPR (GDPR2016a, 11) and the EU AI Act (AIAct24, 10)) have begun to establish binding obligations for organizations to secure sensitive information. The demonstrated harm of breaches in practice, combined with the tightening demands of regulation, points towards an urgent need for robust access control mechanisms that can adapt to the scale and complexity of modern digital ecosystems.
@@
In this paper, we present a system architecture that enables fine-grained access control to protect sensitive information within files in a selective manner. We demonstrate the feasibility of our solution by applying it to visual datasets. Our choice is driven by the fact that such datasets are inherently heterogeneous in nature and often require classification beyond the file level, extending to granularity (such as specific regions, objects, or attributes) of the visual content of individual files (zhao2025visual, 33). Our proposed system utilizes ML techniques to determine where to apply access protection selectively by detecting Privacy Sensitive Objects (PSOs) within the file and classifying these objects based on their varying levels of sensitivity. Regarding how to apply access protection, we make a critical observation that it is commonplace to use blurring- or pixelation-based redaction to protect sensitive information in images. These methods are less secure because redaction only transforms the underlying content into a distorted, lossy representation that is vulnerable to reconstruction attacks, in which an adversary can accurately reconstruct the redacted parts (bland2022story, 4). To address this, our solution employs cryptographic protection that not only provides stronger protection against reconstruction attacks, but also enables controlled reconstruction by design, allowing authorized users to recover protected regions securely through decryption.
Contributions: We present a system that orchestrates ML and cryptographic techniques for fine-grained access control and contributes in the following directions.
@@
PSO Protection: We develop a cryptographic protection component that complements PSO detection through a hybrid encryption scheme and an optimized policy design. The hybrid scheme combines symmetric encryption to encrypt the PSOs and Attribute-Based encryption (ABE) to encrypt the symmetric keys. While symmetric encryption offers performance and scalability, ABE integrates encryption with fine-grained access control within a single mechanism without external enforcement. Also, the policy-driven nature of ABE makes it a suitable candidate for seamless integration with widely deployed access control frameworks (parkinson2022survey, 23). The optimized policy design for ABE reduces operational overhead and enhances scalability by enabling each authorized user to perform a single ABE decryption to efficiently recover all symmetric keys associated with their access rights and by avoiding a naive approach of multiple decryptions for different portions of the data.
@@ 2. Related Work
Computer vision applications heavily employ machine learning methods for three core tasks: object detection, segmentation, and optical character recognition (OCR). Object detection (yolov8, 16, 24) is used to locate, identify, and classify objects within an image, while segmentation (he2017mask, 14, 17, 30) labels each pixel with a pre-defined class. Semantic segmentation labels every pixel in an input image without differentiating between individual objects. In contrast, instance segmentation provides pixel-level labels while also identifying each object separately. State-of-the-art techniques for object detection and segmentation have achieved impressive results in identifying visual components. However, these methods are not equipped to extract text. In contrast, OCR systems (paddleocr, 8, 26) have been developed to recognize text within images. OCRs are widely used to understand scanned documents or identity verification. The combination of these methods are used in various applications, including pose detection, video captioning, and scene graph prediction, and as a fundamental component in self-driving cars.
@@
Orekondy et al. (Orekondy2018Connecting, 21) introduced the first approach for automated redaction of private content from images. They derived a dataset from (orekondy17iccv, 22) by selecting images with privacy-sensitive regions that can be localized for redaction, providing annotations for private objects. They also evaluated semantic segmentation methods for automated redaction via masking. Similarly, Gurari et al. (Gurari2019VizWiz, 13) released a visual privacy recognition dataset, which contains images captured by people with visual impairments, and evaluated unintentional privacy leaks in visual question answering tasks. Building on the notion of visual privacy risk score in (orekondy17iccv, 22), Chen et al. (Chen2021Privattnet, 7) used LSTMs with attention maps to estimate the privacy risk of an entire image. However, their approach does not explicitly localize sensitive objects. More recently, Tay et al. (Tay2024PrivObfNet, 28) combined attention maps with weakly supervised semantic segmentation to predict privacy scores, identify categories of sensitive objects, and generate masks for obfuscation. Although the method in (Tay2024PrivObfNet, 28) improves attention quality compared to previous segmentation models, it focuses primarily on a single object and roughly covers all textual areas while failing to localize text objects. Tseng et al. (Tseng2025BIV, 29) extended the work in (Gurari2019VizWiz, 13) by releasing a new dataset, which contains images with segmented private objects, tailored for the localization of sensitive regions. They also benchmarked several few-shot object detection and segmentation models on this dataset. However, their dataset typically contains only a single annotated privacy sensitive object, reducing the need for fine-grained, multilevel obfuscation mechanisms.
@@
Prior work has explored privacy-preserving data sharing in cloud and IoT environments, including revocable and collaborative ABE schemes (4223236, 3, 2, 31, 32) for dynamic user groups, SeDaSC (Ali2017SeDaSC, 1) which counters insider threats using key shares managed by a trusted third party, and IoT-focused approaches (cryptoeprint:2020/176, 12, 20) such as Symmetric Searchable Encryption for encrypted search and offloading of heavy security operations to edge servers. Other specialized solutions include SecRCNN (Liu2022Privacy, 19), a lightweight privacy-preserving Faster R-CNN framework for medical images using additive secret sharing and edge computing, and a user-centric data space (Robert2025Unlinkable, 25) combining differential privacy with fine-grained access control for unlinkable data sharing via a central intermediary. In contrast, our work embeds fine-grained access control directly into the cryptographic layer via a hybrid Attribute-Based Encryption scheme, efficiently protecting privacy-sensitive objects in visual data without relying on trusted external intermediaries, interactive protocols, or heavy computation on end devices.
@@ 3. Preliminaries
To support fine-grained control, each privacy-sensitive object (PSO) is assigned a sensitivity score on a continuous scale $\alpha,\beta]$, where $\alpha$ denotes the lowest and $\beta$ the highest degree of sensitivity. These scores quantify the relative privacy risk associated with individual PSOs and serve as the basis for grouping and policy assignment. After scoring, POSs are assigned discrete sensitivity groups to simplify policy enforcement. The number of groups $n$ is configurable by the system administrator, organizational policies, privacy regulations, or user-defined requirements. The interval $\alpha,\beta]$ is partitioned into bins $n$, each corresponding to a distinct sensitivity group $G_{1},\dots,G_{n}$. The boundaries of each group $\alpha_{\ell},\beta_{\ell})$, define the sensitivity thresholds. Groups with higher score ranges represent more sensitive information (similar to document classification levels), and therefore require stricter access control and stronger cryptographic protection.
@@
  • $\mathsf{SKE.Dec}(\mathsf{K},c)$: Takes an input a symmetric keu $\mathsf{K}$ and ciphertext $c\in\mathcal{C}$, and ouputs a plaintext $m\in\mathcal{M}$
@@ 4. Proposed Architecture
The Encrypted Data Repository stores the encrypted data and associated metadata provided by the system plane. It is location-agnostic and can reside on the user’s local device, enterprise server, or cloud platform, depending on deployment requirements.
@@ 4.1. AccessPolicy Module
The AccessPolicy module forms the backbone of our system’s privacy and access control framework. It receives the sensitivity group definitions (i.e., the thresholds for each group), the roles defined by the system administrator, and a list of authorized users with their associated attributes. This module is responsible for: This module is responsible for:
@@ 4.2. Detection and Classification Module
Detecting PSOs requires different strategies depending on the modality of the object. Using a single model is often insufficient, since PSOs vary greatly in shape, appearance, and reliance on visual vs. semantic cues. Following the definition in prior work (Orekondy2018Connecting, 21), we categorize PSOs into three groups based on the modality:
@@ 5.1. Training Dataset Preparation
To evaluate the feasibility of our proposed system, we required a dataset that (i) contains multiple PSOs per image to capture contextual relationships, (ii) incorporates some notion of scoring that reflects the perceived sensitivity of the PSOs, and (iii) includes visually localizable annotations so that PSOs can be spatially pinpointed. We investigated four large-scale image privacy datasets that contain annotated PSOs: VISPR (orekondy17iccv, 22), VISPR-Redactions (Orekondy2018Connecting, 21), WizViz (Gurari2019VizWiz, 13), BIV-Priv-Seg (Tseng2025BIV, 29). Table 2 provides a comparison of their properties. While VizWiz and BIV-Priv-Seg costly contain a single PSO per image and lack sensitivity scores, VISPR contains private content that is not visually localizable. VISPR-Redactions satisfied all of our requirements and was therefore selected for evaluating our system.
@@
Annotated Objects Availability Score
VISPR (orekondy17iccv, 22) 5.2 & + &
VISPR-Redactions (Orekondy2018Connecting, 21) 5.6 & & &
WizViz (Gurari2019VizWiz, 13) 1.6 + & +
BIV-Priv-Seg (Tseng2025BIV, 29) 0.9 & & +
@@ 5.2. Algorithm Benchmarking
As stated in Section 4.2, visual PSOs require segmentation to produce precise pixel-level masks to preserve utility. To cover a range of well-known approaches, we fine-tuned four representative segmentation models (with two variants) on the VISPR-Redactions dataset: 1) DeepLabV3+ (chen2018encoder, 6), an encoder-decoder algorithm that leverages atrous (or diluted) convolutions for extracting features at different resolutions, 2) YOLOv8-Seg (s and x variants) (yolov8, 16), an extension of the popular YOLO object detector with mask prediction, 3) SegFormer-B5 (xie2021segformer, 30) which combines hierarchical transformer encoder with a multi layer perceptron decoder, 4) Mask R-CNN (he2017mask, 14) with ResNet-101-FPN and ResNeXt-101-FPN backbones, which extends Faster R-CNN with a parallel mask prediction branch. We measured the performance of segmentation models by calculating the mean Intersection over Union (mIoU). mIoU is defined as the average ratio of intersection over union between predicted and ground-truth masks across all classes.
@@
Textual PSOs (e.g., name, email address, birthdate) require semantic interpretation rather than visual cues. Therefore, as stated in Section 4.2, we first extracted text using OCR and then classified each instance with three transformer-based language models that had fine-tuned over VISPR-Redactions: 1) BERT (devlin2019bert, 9), an encoder-only model that learns contextual word representations, 2) DeBERTa (he2021deberta, 15), which improves BERT with disentangled attention algorithm and mask decoder, 3) MPNet (song2020mpnet, 27), which combines masked and permuted language modeling during training. In addition, we introduced a rule-based correction module (Section 4.3), Post-BERT, which refines the predictions of BERT based on the contextual cues surrounding the detected text. These rule-based adjustments are listed previously in Section 4.3.
@@
Multimodal PSOs (e.g., credit cards, passports, and tickets) in the data set have structured rectangular forms and contain different types of information that must be extracted for accurate identification, thus requiring object detection. We selected four object detection models for their balance in accuracy and speed, and fine-tuned them over VISPR-Redactions: (1) Cascade R-CNN (cai2018cascade, 5), which refines predictions through a sequence of progressively selective detecters against false positives, (2) RetinaNet (lin2017focal, 18), a single-stage detection model designed to handle class imbalance problem with focal loss, (3) Faster R-CNN with ResNeXt-101-FPN backbone (ren2015faster, 24), two-stage detector that combines region proposals with feature pyramids, and (4) YOLOv8 (yolov8, 16), a widely popular single-stage detection model optimized for real-time performance. To further distinguish visually similar objects, we incorporated a Context-Aware Post-Correction (CAPC) module on top of these models: After the detection module, OCR is applied to the bounding box, and the extracted text is used to adjust the predicted label of the identified object, where DeBERTa is used for reclassification. Detection performance is measured by calculating the mean average precision (mAP) when the intersection over union (IoU) threshold varies between $0.5$ and $0.95$, capturing both the localization and the classification accuracy.

### v2

@@ 1. Introduction
The rapid growth of data-driven digital systems and applications has led to the collection and storage of large volumes of sensitive data, ranging from personally identifiable information (PII) to financial transactions and healthcare records. Managing and regulating access to this information has become a critical challenge for individuals and organizations. Beyond protection against external threats, effective access control is a cornerstone of data security, ensuring that only authorized users, applications, and processes can interact with sensitive information. Access control mechanisms regulate and restrict access to resources, systems, or physical areas, typically granting permissions to users based on predefined roles and policies. For example, access to medical records is limited to healthcare professionals, surveillance footage is restricted to security personnel, and digital copies of sealed judicial records are provided only to lawyers and judges. While dealing with such semantically nuanced restrictions, access control mechanisms must operate both reliably and efficiently. The consequences of inadequate access control mechanisms can result in serious financial, reputational, and operational damage as seen in major data breach incidents [^1] https://www.ibm.com/reports/data-breach. Complementing this reality, regulatory frameworks (e.g., the GDPR (European Parliament and Council of the European Union, 2016) and the EU AI Act (European Parliament and Council of the European Union, 2024)) have begun to establish binding obligations for organizations to secure sensitive information. The demonstrated harm of breaches in practice, combined with the tightening demands of regulation, points towards an urgent need for robust access control mechanisms that can adapt to the scale and complexity of modern digital ecosystems.
@@
In this paper, we present a system architecture that enables fine-grained access control to protect sensitive information within files in a selective manner. We demonstrate the feasibility of our solution by applying it to visual datasets. Our choice is driven by the fact that such datasets are inherently heterogeneous in nature and often require classification beyond the file level, extending to granularity (such as specific regions, objects, or attributes) of the visual content of individual files (Zhao et al., 2025). Our proposed system utilizes ML techniques to determine where to apply access protection selectively by detecting Privacy Sensitive Objects (PSOs) within the file and classifying these objects based on their varying levels of sensitivity. Regarding how to apply access protection, we make a critical observation that it is commonplace to use blurring- or pixelation-based redaction to protect sensitive information in images. These methods are less secure because redaction only transforms the underlying content into a distorted, lossy representation that is vulnerable to reconstruction attacks, in which an adversary can accurately reconstruct the redacted parts (Bland et al., 2022). To address this, our solution employs cryptographic protection that not only provides stronger protection against reconstruction attacks, but also enables controlled reconstruction by design, allowing authorized users to recover protected regions securely through decryption.
Contributions: We present a system that orchestrates ML and cryptographic techniques for fine-grained access control, and contribute in the following directions.
@@
PSO Protection: We develop a cryptographic protection component that complements PSO detection through a hybrid encryption scheme and an optimized policy design. The hybrid scheme combines symmetric encryption to encrypt the PSOs and Attribute-Based encryption (ABE) to encrypt the symmetric keys. While symmetric encryption offers performance and scalability, ABE integrates encryption with fine-grained access control within a single mechanism without external enforcement. Also, the policy-driven nature of ABE makes it a suitable candidate for seamless integration with widely deployed access control frameworks (Parkinson and Khan, 2022). The optimized policy design for ABE reduces operational overhead and enhances scalability by enabling each authorized user to perform a single ABE decryption to efficiently recover all symmetric keys associated with their access rights and by avoiding a naive approach of multiple decryptions for different portions of the data.
@@ 2. Related Work
Computer vision applications heavily employ machine learning methods for three core tasks: object detection, segmentation, and optical character recognition (OCR). Object detection (Jocher and others, 2023; Ren et al., 2017) is used to locate, identify, and classify objects within an image, while segmentation (He et al., 2017; Kirillov et al., 2023; Xie et al., 2021) labels each pixel with a pre-defined class. Semantic segmentation labels every pixel in an input image without differentiating between individual objects. In contrast, instance segmentation provides pixel-level labels while also identifying each object separately. State-of-the-art techniques for object detection and segmentation have achieved impressive results in identifying visual components. However, these methods are not equipped to extract text. In contrast, OCR systems (Cui et al., 2025; Smith, 2007) have been developed to recognize text within images. OCRs are widely used to understand scanned documents or identity verification. The combination of these methods are used in various applications, including pose detection, video captioning, and scene graph prediction, and as a fundamental component in self-driving cars.
@@
Orekondy et al. (Orekondy et al., 2018) introduced the first approach for automated redaction of private content from images. They derived a dataset from (Orekondy et al., 2017) by selecting images with privacy-sensitive regions that can be localized for redaction, providing annotations for private objects. They also evaluated semantic segmentation methods for automated redaction via masking. Similarly, Gurari et al. (Gurari et al., 2019) released a visual privacy recognition dataset, which contains images captured by people with visual impairments, and evaluated unintentional privacy leaks in visual question answering tasks. Building on the notion of visual privacy risk score in (Orekondy et al., 2017), Chen et al. (Chen et al., 2021) used LSTMs with attention maps to estimate the privacy risk of an entire image. However, their approach does not explicitly localize sensitive objects. More recently, Tay et al. (Tay et al., 2024) combined attention maps with weakly supervised semantic segmentation to predict privacy scores, identify categories of sensitive objects, and generate masks for obfuscation. Although the method in (Tay et al., 2024) improves attention quality compared to previous segmentation models, it focuses primarily on a single object and roughly covers all textual areas while failing to localize text objects. Tseng et al. (Tseng et al., 2025) extended the work in (Gurari et al., 2019) by releasing a new dataset, which contains images with segmented private objects, tailored for the localization of sensitive regions. They also benchmarked several few-shot object detection and segmentation models on this dataset. However, their dataset typically contains only a single annotated privacy sensitive object, reducing the need for fine-grained, multilevel obfuscation mechanisms.
@@
Prior work has explored privacy-preserving data sharing in cloud and IoT environments, including revocable and collaborative ABE schemes (Bethencourt et al., 2007; Bakas and Michalas, 2019; Xu et al., 2018; Xue et al., 2019) for dynamic user groups, SeDaSC (Ali et al., 2017) which counters insider threats using key shares managed by a trusted third party, and IoT-focused approaches (Frimpong et al., 2020; Mollah et al., 2017) such as Symmetric Searchable Encryption for encrypted search and offloading of heavy security operations to edge servers. Other specialized solutions include SecRCNN (Liu et al., 2022), a lightweight privacy-preserving Faster R-CNN framework for medical images using additive secret sharing and edge computing, and a user-centric data space (Röbert et al., 2025) combining differential privacy with fine-grained access control for unlinkable data sharing via a central intermediary. In contrast, our work embeds fine-grained access control directly into the cryptographic layer via a hybrid Attribute-Based Encryption scheme, efficiently protecting privacy-sensitive objects in visual data without relying on trusted external intermediaries, interactive protocols, or heavy computation on end devices.
@@ 3. Preliminaries
To support fine-grained control, each privacy-sensitive object (PSO) is assigned a sensitivity score on a continuous scale $\alpha,\beta]$, where $\alpha$ denotes the lowest and $\beta$ the highest degree of sensitivity. These scores quantify the relative privacy risk associated with individual PSOs and serve as the basis for grouping and policy assignment. After scoring, PSOs are assigned discrete sensitivity groups to simplify policy enforcement. The number of groups $n$ is configurable by the system administrator, organizational policies, privacy regulations, or user-defined requirements. The interval $\alpha,\beta]$ is partitioned into bins $n$, each corresponding to a distinct sensitivity group $G_{1},\dots,G_{n}$. The boundaries of each group $\alpha_{\ell},\beta_{\ell})$, define the sensitivity thresholds. Groups with higher score ranges represent more sensitive information (similar to document classification levels), and therefore require stricter access control and stronger cryptographic protection.
@@
  • $\mathsf{SKE.Dec}(\mathsf{K},c)$: Takes an input a symmetric key $\mathsf{K}$ and ciphertext $c\in\mathcal{C}$, and outputs a plaintext $m\in\mathcal{M}$
@@ 4. Proposed Architecture
The Encrypted Data Repository stores the encrypted data and associated metadata provided by the system plane. It is location-agnostic and can reside on the enterprise server or cloud platform, depending on the deployment requirements. Users can access encrypted files, but they require the correct keys to decrypt them.
@@ 4.1. AccessPolicy Module
The AccessPolicy module forms the backbone of our system’s privacy and access control framework. It receives the sensitivity group definitions (i.e., the thresholds for each group), the roles defined by the system administrator, and a list of authorized users with their associated attributes. This module is responsible for:
@@ 4.2. Detection and Classification Module
Detecting PSOs requires different strategies depending on the modality of the object. Using a single model is often insufficient, since PSOs vary greatly in shape, appearance, and reliance on visual vs. semantic cues. Following the definition in prior work (Orekondy et al., 2018), we categorize PSOs into three groups based on the modality:
@@ 5.1. Training Dataset Preparation
To evaluate the feasibility of our proposed system, we required a dataset that (i) contains multiple PSOs per image to capture contextual relationships, (ii) incorporates some notion of scoring that reflects the perceived sensitivity of the PSOs, and (iii) includes visually localizable annotations so that PSOs can be spatially pinpointed. We investigated four large-scale image privacy datasets that contain annotated PSOs: VISPR (Orekondy et al., 2017), VISPR-Redactions (Orekondy et al., 2018), WizViz (Gurari et al., 2019), BIV-Priv-Seg (Tseng et al., 2025). Table 2 provides a comparison of their properties. While VizWiz and BIV-Priv-Seg costly contain a single PSO per image and lack sensitivity scores, VISPR contains private content that is not visually localizable. VISPR-Redactions satisfied all of our requirements and was therefore selected for evaluating our system.
@@
Annotated PSOs Availability Score
VISPR (Orekondy et al., 2017) 5.2
VISPR-Redactions (Orekondy et al., 2018) 5.6
WizViz (Gurari et al., 2019) 1.6
BIV-Priv-Seg (Tseng et al., 2025) 0.9
@@ 5.2. Algorithm Benchmarking
As stated in Section 4.2, visual PSOs require segmentation to produce precise pixel-level masks to preserve utility. To cover a range of well-known approaches, we fine-tuned four representative segmentation models (with two variants) on the VISPR-Redactions dataset: 1) DeepLabV3+ (Chen et al., 2018), an encoder-decoder algorithm that leverages atrous (or diluted) convolutions for extracting features at different resolutions, 2) YOLOv8-Seg (s and x variants) (Jocher and others, 2023), an extension of the popular YOLO object detector with mask prediction, 3) SegFormer-B5 (Xie et al., 2021) which combines hierarchical transformer encoder with a multi layer perceptron decoder, 4) Mask R-CNN (He et al., 2017) with ResNet-101-FPN and ResNeXt-101-FPN backbones, which extends Faster R-CNN with a parallel mask prediction branch. We measured the performance of segmentation models by calculating the mean Intersection over Union (mIoU). mIoU is defined as the average ratio of intersection over union between predicted and ground-truth masks across all classes.
@@
Textual PSOs (e.g., name, email address, birthdate) require semantic interpretation rather than visual cues. Therefore, as stated in Section 4.2, we first extracted text using OCR and then classified each instance with three transformer-based language models that had fine-tuned over VISPR-Redactions: 1) BERT (Devlin et al., 2019), an encoder-only model that learns contextual word representations, 2) DeBERTa (He et al., 2021), which improves BERT with disentangled attention algorithm and mask decoder, 3) MPNet (Song et al., 2020), which combines masked and permuted language modeling during training. In addition, we introduced a rule-based correction module (Section 4.3), Post-BERT, which refines the predictions of BERT based on the contextual cues surrounding the detected text. These rule-based adjustments are listed previously in Section 4.3.
@@
Multimodal PSOs (e.g., credit cards, passports, and tickets) in the data set have structured rectangular forms and contain different types of information that must be extracted for accurate identification, thus requiring object detection. We selected four object detection models for their balance in accuracy and speed, and fine-tuned them over VISPR-Redactions: (1) Cascade R-CNN (Cai and Vasconcelos, 2018), which refines predictions through a sequence of progressively selective detecters against false positives, (2) RetinaNet (Lin et al., 2017), a single-stage detection model designed to handle class imbalance problem with focal loss, (3) Faster R-CNN with ResNeXt-101-FPN backbone (Ren et al., 2017), two-stage detector that combines region proposals with feature pyramids, and (4) YOLOv8 (Jocher and others, 2023), a widely popular single-stage detection model optimized for real-time performance. To further distinguish visually similar objects, we incorporated a Context-Aware Post-Correction (CAPC) module on top of these models: After the detection module, OCR is applied to the bounding box, and the extracted text is used to adjust the predicted label of the identified object, where DeBERTa is used for reclassification. Detection performance is measured by calculating the mean average precision (mAP) when the intersection over union (IoU) threshold varies between $0.5$ and $0.95$, capturing both the localization and the classification accuracy.